OpenStudy (anonymous):

What is 0.6(y + 3) = 4.8

OpenStudy (lgbasallote):

divide both sides by 0.6 you'll get y + 3 = 8 subtract 3 from both sides y = 8 - 3 y = 5

OpenStudy (anonymous):

y=5